OpenStudy (anonymous):

A sequence has a third term of 18, and a common ratio of 3. What is the fifth term? 6 54 162 486

OpenStudy (ybarrap):

$$ a_n = a\,r^{n-1}. $$ where r is your common ratio and n denotes the term number, 1,2,3, etc $$ a_5=6\times3^{5-1}=? $$

OpenStudy (ybarrap):

$$ a_0=6\times3^{-1}=6/3=2\\ a_1=6\\ a_2=6\times3=18\\ a_3=6\times3^2=54\\ a_4=6\times3^3=162\\ $$ These are called geometric series: http://en.wikipedia.org/wiki/Geometric_progression

OpenStudy (triciaal):

a, b, c given c = 18 ratio = 3 c/b =3

OpenStudy (triciaal):

18 x 3 = 54 54 x 3 = 162